# Antiferromagnetism and Tightly Bound Cooper Pairs Induced by Kinetic Frustration

### v1

@@ A.2 Antiferromagnetic order through inter-sublattice hopping $t_{1}$
In the ferromagnetic configuration ($\hat{n}_{1}=\hat{n}_{2}=\hat{z}$), where all spins are uniformly polarized, the $t_{1}$ hopping processes between sublattices interfere destructively, causing the probability amplitude for a hole to hop to the opposite sublattice to cancel exactly. As illustrated in Fig. S1(a), when a hole attempts to move across sublattices, the resulting amplitudes acquire opposite signs due to the phase factor $(-1)^{(x_{1}+x_{2}-1)}$ in the wavefunction. This destructive interference ensures that the original $t_{2}$-only ground state remains an exact eigenstate, even in the presence of a finite $t_{1}$, as long as the system remains fully spin-polarized.
However, when the two sublattices are polarized in different spin directions, the spin mismatch left behind by a hole’s motion differs between the various $t_{1}$ hopping paths. This asymmetry removes the exact cancellation described previously and enables the system to lower its energy via second-order processes—namely, a hole hops to the opposite sublattice and then returns. These second-order virtual hoppings respect the underlying $SU(2)$ symmetry and couple only sublattice spin states with magnetic quantum numbers $m$ and $m\pm 1$. As a result, the corresponding energy reduction must take the following form:
$\Delta E=c(\mathbf{S}_{1}\cdot\mathbf{S}_{2}-\mathbf{S}_{1}^{2})$ (6)
Here, $\mathbf{S}_{1}$ and $\mathbf{S}2$ denote the total spins of the two sublattices, with equal magnitude, and $c>0$ is a positive constant. This implies that the lowest-energy state resides in the $S{\text{total}}=0$ sector, where $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are aligned antiparallel to form a singlet.
The Néel order correlation function—also referred to as the square of the staggered magnetization—serves as the order parameter for this phase and is defined as follows:
$m(\bm{q})^{2}=\frac{1}{N_{\mathrm{s}}^{2}}\sum_{i,j}e^{i\bm{q}\cdot\left(\bm{r}_{i}-\bm{r}_{j}\right)}\left\langle\Psi_{G}|\bm{S}_{i}\cdot\bm{S}_{j}|\Psi_{G}\right\rangle$ (7)
with $q=(\pi,\pi)$. Here, $N_{\mathrm{s}}$ is the total number of sites in the system. For the ground state singlet wavefunction in $S{\text{total}}=0$ sector, the Néel order parameter is: $1/4-1/N_{\mathrm{s}}^{2}$. In the thermodynamic limit, this approaches 0.25.
@@ A.3 Robustness of the kinetically stabilized Néel order for finite $t_{1}/t_{2}$
On top of the perturbative wavefunction approach, we now verify the stability of the kinetically stabilized Néel order at large $t_{1}$ regime through numerical exact diagonalization using the XDiag and QuSpin packages [52, 53, 54]. Fig. S1(b) shows the ground state remains in the $S_{\text{total}}=0$ sector even as $t_{2}/t_{1}$ down to values around 0.5. The Néel order parameter there maintains values close to the maximum 0.25 from $t_{2}/t_{1}\to\infty$ down to $t_{2}/t_{1}\sim 0.5$ [Fig. S1(c)].
@@
Figure S1: (a) The schematic diagram for the cancellation of hopping processes between opposite sublattices. The phase of superposition coefficient for $|\Psi_{m_{1}=m_{2}=J}\rangle$ is marked on the graph. (b) The spin gap as energy difference between ground state energy at different total spin sector, calculated in a $4\times 4$ PBC cluster. (c) The Néel order parameter: $m(\bm{q})^{2}=\frac{1}{N_{\mathrm{s}}^{2}}\sum_{i,j}e^{i\bm{q}\cdot\left(\bm{r}_{i}-\bm{r}_{j}\right)}\left\langle\bm{S}_{i}\cdot\bm{S}_{j}\right\rangle$ with respect to $t_{2}/t_{1}$, calculated with exact diagonalization on a $4\times 4$ and $4\times 6$ PBC cluster.
Appendix B $t-J$ model in the exact solvable $\lambda=\Delta=0$ limit
@@ B.1 Hamiltonian and wavefunction structure
Building on our previous kinetic magnetic order analysis in the infinite-$U$ limit with small $\lambda$, we now turn to the exactly solvable limit at $\lambda=\Delta=0$ with finite $U$. In this regime, the nearest-neighbor hopping $t_{1}$ vanishes, and the system features Ising-type spin interactions.
The Hamiltonian simplifies to:
$H$ $=$ $-t_{2}\sum_{\langle\langle jl\rangle\rangle,\sigma}c_{j\sigma}^{\dagger}c_{l\sigma}+h.c.+\sum_{\langle jl\rangle}\frac{J_{1}}{2}\left[S^{z}_{j}S^{z}_{l}-\frac{n_{j}n_{l}}{4}\right]$ (8)
$+\sum_{\langle\langle jl\rangle\rangle}\frac{J_{2}}{2}\left[S^{z}_{j}S^{z}_{l}-\frac{n_{j}n_{l}}{4}\right]$
@@
The hole-hole interaction arises solely from differences in Ising interaction energies between adjacent and isolated hole configurations. As detailed in the main text, when two holes occupy neighboring lattice sites, the number of broken antiferromagnetic bonds is reduced from eight to seven. This reduction effectively generates an attractive interaction of magnitude $J_{1}/2$ between vertically or horizontally adjacent holes.
@@
$|\Psi_{G}\rangle=\sum_{r_{1}\in A}\sum_{r_{2}\in B}\psi(r_{1},r_{2})c_{1,\uparrow}(r_{1})c_{2,\downarrow}(r_{2})|\uparrow\downarrow...\uparrow\downarrow\rangle$ (9)
@@
For translationally invariant systems, we can fix the center-of-mass momentum to $G$ and rewrite the wavefunction in terms of the relative coordinate $R=r_{2}-r_{1}$:
$\psi(r_{1},r_{2})=\tilde{\psi}(R)e^{iGr_{1}}$ (10)
The relative wavefunction $\tilde{\psi}(R)$ satisfies the equation:
$\sum_{i}t_{ij}(1+e^{iG(R_{i}-R_{j})})\tilde{\psi}(R_{i})-\frac{J_{1}}{2}\delta(|R_{i}|-a)\tilde{\psi}(R_{j})=E\tilde{\psi}(R_{j})$ (11)
@@ B.2 Pairing symmetry and Cooper pair wavefunction
Analyzing the two-hole Cooper pair wavefunction yields essential insights into the pairing symmetry and spatial structure in the $(\lambda,\Delta)=(0,0)$ limit. By solving the eigenvalue equation (Eq. 11) for various center-of-mass momenta $G$, we can clearly characterize the underlying pairing mechanism in this exact solvable limit.
Our analysis confirms that the minimum energy occurs at $G=(0,0)$, where the Hamiltonian becomes real and time-reversal invariant. In this momentum sector, the system naturally selects $d_{x^{2}-y^{2}}$-wave pairing symmetry. This symmetry emerges from the alternating signs of the wavefunction across the lattice, minimizing kinetic energy governed by $t_{2}$.
Fig. S2 illustrates how the Cooper pair wavefunction $\tilde{\psi}(R)$ changes with interaction strength. Panel (a) shows the wavefunction at $U=10t_{1}$ (weak exchange with $J_{1}\approx 0.4t_{1}$), while panel (b) shows it at $U=2t_{1}$ (strong exchange with $J_{1}\approx 2t_{1}$). The $d_{x^{2}-y^{2}}$ symmetry is evident from the sign changes (color transitions between red and blue) along different directions.
The spatial extent of the Cooper pair varies dramatically with interaction strength. At large $U$ (panel a), the slow spatial decay of the wavefunction indicates a weakly bound state characterized by a long coherence length $\xi\gg a$. Conversely, at smaller $U$ (panel b), the wavefunction decays rapidly with distance, revealing a tightly bound pair with a coherence length $\xi$ of only a few lattice spacings.
@@ B.3 Continuum limit analysis and binding energy
$\tilde{\psi}(R_{i})=(-1)^{(X_{i}-1)}\sqrt{2}a\tilde{\phi}(R_{i})$ (12)
@@
[…]
@@
$-\frac{\hbar^{2}}{2m^{*}}\nabla^{2}\tilde{\phi}(R)-v_{0}\delta(r)\tilde{\phi}(R)=(E+8t_{2})\tilde{\phi}(R)$ (13)
@@
The bound-state wavefunction correspondingly takes the form at $|R|\gg 1$:
$\tilde{\phi}(R)\sim K_{0}\left(\frac{\sqrt{2m^{*}\Delta_{g}}}{\hbar}|R|\right)$ (15)
where $K_{0}$ is the modified Bessel function of the second kind of order zero.
For distances where $\frac{\sqrt{2m^{*}\Delta_{g}}}{\hbar}R\gg 1$, the wavefunction assumes the asymptotic form:
$\tilde{\phi}(R)\sim\frac{1}{\sqrt{R}}e^{-\frac{\sqrt{2m^{*}\Delta_{g}}}{\hbar}R}$ (16)
This expression directly connects the binding energy $\Delta_{g}$ to the Cooper pair coherence length $\xi=\frac{\hbar}{\sqrt{2m^{*}\Delta_{g}}}\sim e^{\frac{2\pi|t_{2}|}{J_{1}}}$. At small $J_{1}/t_{2}$ values, this length grows exponentially, explaining the nearly uniform amplitude observed in Fig. S2(a).
Figure S2: (a) The wavefunction $\tilde{\psi}(R)$ of the Cooper pair under parameter $(\lambda,\Delta)=(0,0)$ and $U=10t_{1}$. Black dot marks the origin. A clear $d_{x^{2}-y^{2}}$ wave symmetry can be seen. (b) The wavefunction with same parameter but at $U=2t_{1}$. (c) The scaling of $\log(r|\tilde{\psi}(r)|^{2})$ with respect to $r$. Linear dependency is observed at large $r$ but still away from boundary. Smaller system size tends to give incorrect scaling.
@@ B.4 Finite size effects
Two key observables show pronounced finite-size sensitivity:
First, the Cooper pair wavefunction’s spatial structure changes dramatically. Fig. S2(c) demonstrates how PBC artificially enhance hole-hole correlations, especially in small systems where holes can interact with their periodic images. This enhancement can be severe enough to reverse the slope when fitting correlation functions to Eq. 16, misleadingly suggesting repulsion rather than attraction.
Second, the binding energy calculation requires careful consideration of finite-size effects. The conventional binding energy definition $\Delta_{g}(k=(0,0))=E_{1}(k=(0,0))-E_{0}(k=(0,0))$ contains artificial finite-size contributions due to the discrete momentum grid, which creates a spurious gap:
$\Delta_{g}(k=(0,0))_{\text{noninteracting}}=8t_{2}\left(1-\cos\left(\frac{2\pi}{L}\right)\right).$ (17)
Hereafter for convenient, we assume the lattice to be square: $L_{x}=L_{y}=L$. This artificial contribution makes measuring the true interaction-induced binding difficult, especially when genuine binding is weak: $\Delta_{g}\ll\Delta_{g}(k=(0,0))_{\text{noninteracting}}$.
We solve this problem by calculating the gap at shifted momentum $k=(0,2\pi/L)$, which eliminates the non-interacting gap. This momentum shift provides an additional benefit: it creates mixed boundary conditions (periodic along $x$, antiperiodic along $y$) for the relative wavefunction in Eq. 11. These boundary conditions generate bonding and antibonding states between two holes, with the antibonding state effectively making holes blind to the attraction when two are near and producing a continuum-like wavefunction (See Fig. S7(d)). The energy splitting between these states accurately reflects if the hole-hole interaction is attractive or repulsive.
Fig. S3 demonstrates how our shifted-momentum approach improves convergence with system size. The shifted-momentum gap converges much faster than the conventional definition, allowing reliable extraction of pairing strength even from modest system sizes. We apply this approach consistently throughout our study of fluctuation effects on the binding gap.
Figure S3: Comparison of binding energy calculations using different momentum sectors and system sizes. (a) The binding gap $\Delta_{g}(k)=E_{1}(k)-E_{0}(k)$ measured at $k=(0,0)$ for various system sizes and interaction strengths ($U/t_{1}=1-5$). (b) The improved binding gap measured at $k=(0,2\pi/L)$, showing much faster convergence with system size.
Appendix C The Hilbert space truncation in ED

### v2

@@ A.2 Antiferromagnetic order through inter-sublattice hopping $t_{1}$
In the ferromagnetic configuration ($\hat{\bm{n}}_{1}=\hat{\bm{n}}_{2}=\hat{\bm{z}}$), where all spins are uniformly polarized, the $t_{1}$ hopping processes between sublattices interfere destructively, causing the probability amplitude for a hole to hop to the opposite sublattice to cancel exactly. As illustrated in Fig. S1(a), when a hole attempts to move across sublattices, the resulting amplitudes acquire opposite signs due to the phase factor $(-1)^{(x_{1}+x_{2}-1)}$ in the wavefunction. This destructive interference ensures that the original $t_{2}$-only ground state remains an exact eigenstate, even in the presence of a finite $t_{1}$, as long as the system remains fully spin-polarized.
However, when the two sublattices are polarized in opposite spin directions, the spin mismatch left behind by a hole’s motion differs between the various $t_{1}$ hopping paths. This asymmetry removes the exact cancellation described previously and enables the system to lower its energy via second-order processes—namely, a hole hops to the opposite sublattice and then returns. These second-order virtual hoppings respect the underlying $SU(2)$ symmetry and couple only sublattice spin states with magnetic quantum numbers $m$ and $m\pm 1$. As a result, the corresponding energy reduction takes the following form:
$\Delta E=c(\bm{S}_{1}\cdot\bm{S}_{2}-\bm{S}_{1}^{2})$ (6)
Here, $\bm{S}_{1}$ and $\bm{S}_{2}$ denote the total spins of the two sublattices, with equal magnitude, and $c>0$ is a positive constant. This implies that the lowest-energy state resides in the $S_{\text{total}}=0$ sector, where $\bm{S}_{1}$ and $\bm{S}_{2}$ are aligned antiparallel to form a singlet ground state.
The Néel order parameter or staggered magnetization that emerges in the thermodynamic limit can be extrapolated from a finite size scaling study of the square of the staggered magnetization normalized by the total number of sties $N_{\mathrm{s}}$:
$m^{2}(\bm{q})=\frac{1}{N_{\mathrm{s}}^{2}}\sum_{i,j}e^{i\bm{q}\cdot\left(\bm{r}_{i}-\bm{r}_{j}\right)}\left\langle\Psi_{G}|\bm{S}_{i}\cdot\bm{S}_{j}|\Psi_{G}\right\rangle$ (7)
with $\bm{q}=(\pi,\pi)$. For the ground state singlet wavefunction in $S_{\text{total}}=0$ sector, we have $m^{2}(\bm{q})=(1/4-1/N^{2}_{\mathrm{s}})$, which approaches 0.25 in the thermodynamic limit.
@@ A.3 Robustness of the kinetically stabilized Néel order for finite $t_{1}/t_{2}$
On top of the perturbative wavefunction approach, we now verify the stability of the kinetically stabilized Néel order at large $t_{1}$ regime through numerical exact diagonalization using the XDiag and QuSpin packages [56, 57, 58]. Fig. S1(b) shows the ground state remains in the $S_{\text{total}}=0$ sector even as $t_{2}/t_{1}$ down to values around 0.5. The Néel order parameter there maintains values close to the maximum 0.25 from $t_{2}/t_{1}\to\infty$ down to $t_{2}/t_{1}\sim 0.5$ [Fig. S1(c)]. This result is consistent with large Néel order parameter when a single hole is doped under intermediate $t_{2}/t_{1}$ [27].
@@
Figure S1: (a) Schematic diagram for the cancellation of hopping processes between opposite sublattices. The phase of superposition coefficient for $|\Psi_{m_{1}=m_{2}=J}\rangle$ is marked on the graph. (b) Spin gaps for each total spin sector calculated in a $4\times 4$ PBC cluster. (c) Square of the Néel order parameter: $m(\bm{q})^{2}=\frac{1}{N_{\mathrm{s}}^{2}}\sum_{i,j}e^{i\bm{q}\cdot\left(\bm{r}_{i}-\bm{r}_{j}\right)}\left\langle\bm{S}_{i}\cdot\bm{S}_{j}\right\rangle$ with respect to $t_{2}/t_{1}$ calculated from ED of $4\times 4$ and $4\times 6$ clusters with PBC.
Appendix B Exactly solvable limit $\lambda=\Delta=0$
@@ B.1 Hamiltonian and wavefunction structure
Building on our previous kinetic magnetic order analysis in the infinite-$U$ limit with small $\lambda$, we now turn to the exactly solvable limit $\lambda=\Delta=0$. In this regime, the nearest-neighbor hopping $t_{1}$ vanishes, and the Hamiltonian features Ising-type spin interactions:
$H$ $=$ $-t_{2}\sum_{\langle\langle jl\rangle\rangle,\sigma}c_{j\sigma}^{\dagger}c_{l\sigma}+h.c.+\frac{J_{1}}{2}\sum_{\langle jl\rangle}\left[S^{z}_{j}S^{z}_{l}-\frac{n_{j}n_{l}}{4}\right]$ (8)
$+\frac{J_{2}}{2}\sum_{\langle\langle jl\rangle\rangle}\left[S^{z}_{j}S^{z}_{l}-\frac{n_{j}n_{l}}{4}\right],$
@@
The hole-hole interaction arises solely from differences in Ising interaction energies between adjacent and isolated hole configurations. As detailed in the main text, when two holes occupy neighboring lattice sites, the number of broken antiferromagnetic bonds is reduced from eight to seven. This reduction effectively generates an attractive interaction of magnitude $J_{1}/2$ between adjacent holes.
@@
$|\Psi_{G}\rangle=\sum_{\bm{r}_{1}\in A,\bm{r}_{2}\in B}\psi(\bm{r}_{1},\bm{r}_{2})c_{1\uparrow}c_{2\downarrow}|\uparrow\downarrow...\uparrow\downarrow\rangle$ (9)
@@
For translationally invariant systems, we can fix the center-of-mass momentum to $\bm{K}$ and rewrite the wavefunction in terms of the center-of-mass coordinate $\bm{R}=\frac{\bm{r}_{1}+\bm{r}_{2}}{2}$ and relative coordinate $\bm{\rho}=\bm{r}_{1}-\bm{r}_{2}$:
$\psi(\bm{r}_{1},\bm{r}_{2})=\tilde{\psi}_{\bm{\rho}}e^{i\bm{K}\cdot\bm{R}}$ (10)
The relative wavefunction $\tilde{\psi}_{\bm{\rho}}$ satisfies the equation:
$\sum_{j}t_{jl}\cos\left(\frac{1}{2}\bm{K}\cdot(\bm{\rho}_{j}-\bm{\rho}_{l})\right)\tilde{\psi}_{\bm{\rho}_{j}}-\frac{J_{1}}{2}\delta(|\bm{\rho}_{j}|-a)\tilde{\psi}_{\bm{\rho}_{l}}=E\tilde{\psi}_{\bm{\rho}_{l}}$ (11)
@@ B.2 Pairing symmetry and Cooper pair wavefunction
Analyzing the two-hole Cooper pair wavefunction yields essential insights into the pairing symmetry and spatial structure in the $(\lambda,\Delta)=(0,0)$ limit. By solving the eigenvalue equation Eq. (11) for various center-of-mass momenta $\bm{K}$, we can clearly characterize the underlying pairing mechanism in this exact solvable limit.
Our analysis confirms that the global energy minimum occurs in $\bm{K}=(0,0)$ sector, where the Hamiltonian is real and time-reversal invariant. The corresponding bound state has $d_{x^{2}-y^{2}}$-wave pairing symmetry emerging from the positive tunneling $t_{2}$ which leads to a sign alternating sign alternation of one hole moving around the other one. In other words, the $d_{x^{2}-y^{2}}$ symmetry of the pair is a direct consequence of the kinetic energy frustration.
Fig. S2 illustrates how the Cooper pair wavefunction $\tilde{\psi}_{\bm{\rho}}$ changes with the interaction strength $U$. Here we include values of $U$ for which the original $t-J$ model is obviously invalid, as a way to tune the effective NN attractive interaction estimating the magnitude of the attraction that would be required to obtain a coherence length $\xi$ of order a few lattice spaces. Panel (a) shows the wavefunction for $U=10t_{1}$ (weak exchange with $J_{1}\approx 0.4t_{1}$), while panel (b) shows it for $U=2t_{1}$ (strong exchange with $J_{1}\approx 2t_{1}$). The $d_{x^{2}-y^{2}}$ symmetry is evident from the sign changes (color transitions between red and blue) along different directions.
As expected, $\xi$ varies dramatically with interaction strength. For large $U$ (panel a), the slow spatial decay of the wavefunction indicates a weakly bound state characterized by a coherence length $\xi$ several orders of magnitude higher than $a$. Conversely, at smaller $U$ (panel b), the wavefunction decays rapidly with distance, revealing a tightly bound pair with a coherence length $\xi$ of only a few lattice spacings.
@@ B.3 Continuum limit analysis and binding energy
$\tilde{\psi}_{\bm{\rho}_{i}}=(-1)^{(X_{i}-1)}\sqrt{2}a\tilde{\phi}_{\bm{\rho}_{i}}$ (12)
@@
The resulting effective Hamiltonian in the continuum limit comprises two main components: a kinetic term from $t_{2}$ hopping: $-8t_{2}-4t_{2}a^{2}\int\tilde{\phi}(\bm{\rho})^{*}\nabla^{2}\tilde{\phi}_{\bm{\rho}}\,d\rho^{2}$ and a potential term originating from the nearest-neighbor Ising interaction: $v_{0}|\tilde{\phi}(0)|^{2}$ where $v_{0}=4a^{2}J_{1}$
@@
$-\frac{\hbar^{2}}{2m^{*}}\nabla^{2}\tilde{\phi}_{\bm{\rho}}-v_{0}\delta(\bm{\rho})\tilde{\phi}_{\bm{\rho}}=(E+8t_{2})\tilde{\phi}_{\bm{\rho}}$ (13)
@@
The bound-state wavefunction correspondingly takes the form at $|\rho|\gg 1$:
$\tilde{\phi}_{\bm{\rho}}\sim K_{0}\left(\frac{\sqrt{2m^{*}\Delta_{g}}}{\hbar}\rho\right)$ (15)
where $K_{0}$ is the modified Bessel function of the second kind of order zero and $\rho\equiv|\bm{\rho}|$.
For distances where $\frac{\sqrt{2m^{*}\Delta_{g}}}{\hbar}\rho\gg 1$, the wavefunction assumes the asymptotic form:
$\tilde{\phi}_{\bm{\rho}}\sim\frac{1}{\sqrt{\rho}}e^{-\frac{\sqrt{2m^{*}\Delta_{g}}}{\hbar}\rho}$ (16)
This expression directly connects the binding energy $\Delta_{g}$ to the Cooper pair coherence length
$\xi=\frac{\hbar}{\sqrt{2m^{*}\Delta_{g}}}\sim e^{\frac{2\pi|t_{2}|}{J_{1}}}.$ (17)
$\xi$ grows exponentially in $|t_{2}|/J_{1}$, explaining the nearly uniform amplitude observed in Fig. S2(a).
Figure S2: (a) Wavefunction $\tilde{\psi}_{\bm{\rho}}$ of the relative coordinate of the Cooper pair for $U=10t_{1}$ and $\lambda=\Delta=0$. The black dot marks the origin. A clear $d_{x^{2}-y^{2}}$-wave symmetry is observed. (b) Same as in (a), but for $U=2t_{1}$. (c) Scaling of $\log\left[\rho|\tilde{\psi}_{\bm{\rho}}|^{2}\right]$ as a function of $\rho$. A linear dependence is seen at large distances, away from the system boundaries. Smaller system sizes tend to yield incorrect scaling behavior.
@@ B.4 Finite size effects
Two key observables exhibit pronounced finite-size sensitivity.
First, the spatial structure of the Cooper pair wavefunction is strongly affected. Fig. S2(c) shows that periodic boundary conditions (PBC) artificially enhance hole-hole correlations, particularly in small systems where holes can interact with their periodic images. This artifact can be severe enough to reverse the sign of the slope when fitting correlation functions to Eq. (16), leading to the misleading appearance of repulsion rather than attraction.
Second, the calculation of the binding energy requires careful treatment of finite-size effects. The conventional definition, $\Delta_{g}[\bm{k}=(0,0)]=E_{1}[\bm{k}=(0,0)]-E_{0}[\bm{k}=(0,0)]$, includes spurious finite-size contributions arising from the discreteness of the momentum grid. For non-interacting particles, this results in an artificial finite size gap:
$\Delta_{g}[\bm{k}=(0,0)]_{\text{free}}=8t_{2}\left[1-\cos\left(\frac{2\pi}{L}\right)\right].$ (18)
For convenience, we assume a square lattice with $L_{x}=L_{y}=L$. This artificial contribution complicates the detection of true interaction-induced binding, especially in the regime of weak pairing, where $\Delta_{g}\ll\Delta_{g}[\bm{k}=(0,0)]_{\text{free}}$ or $L<\xi$.
To mitigate this problem we calculate the gap at shifted momentum $\bm{k}=(0,2\pi/L)$, which eliminates the non-interacting gap. This momentum shift provides an additional benefit: it creates mixed boundary conditions (periodic along $x$, antiperiodic along $y$) for the relative wavefunction in Eq. 11. These boundary conditions generate bonding and antibonding states between two holes, with the antibonding state effectively making holes blind to the attraction and producing a continuum-like wavefunction (See Fig. S7(d)). The energy splitting between these states accurately reflects if the hole-hole interaction is attractive or repulsive.
Fig. S3 demonstrates how our shifted-momentum approach improves convergence with system size. The shifted-momentum gap converges faster than the conventional definition, allowing a more reliable extraction of pairing strength even from modest system sizes. We apply this approach consistently throughout our study of spin fluctuation effects on the binding gap.
Figure S3: Comparison of binding energy calculations using different momentum sectors and system sizes. (a) Binding gap $\Delta_{g}(\bm{k})=E_{1}(\bm{k})-E_{0}(\bm{k})$ measured at $\bm{k}=(0,0)$ for various system sizes and interaction strengths ($U/t_{1}=1-5$). (b) Numerically improved binding gap measured at $\bm{k}=(0,2\pi/L)$, showing faster convergence with system size.
Appendix C Hilbert space truncation